OpenStudy (anonymous):

jim has a bag that contained 5 blue 5 red and 5 green marbles. He drew out 1 marble at a time and then placed it back into the bag. what is the theoretical probability of drawing a blue marble a red marble an then a green marble.

OpenStudy (anonymous):

what is the theoretical probability that 

OpenStudy (anonymous):

i thought it was WITH replacement

OpenStudy (anonymous):

what is the experimental probability of drawing a blue marble then a red marble and then a green marble what is the experimental probability that jim does not draw a red marble

OpenStudy (mathmate):

Number of red marbles = number of green marbles = number of blue marbles = 15 Total number of marbles = 15 Draws with replacement P(R)=P(G)=P(B)=\(\frac1 3\) To get BRG in order is a three step experiments in which all steps are independent P(BRG in order)=P(B)*P(R)*P(G)=\(\frac1 3\)*\(\frac1 3\)*\(\frac1 3\)=1/27 Not sure about the question following your initial post.